40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9484     Entry time: Wed Dec 18 00:26:15 2013     In reply to: 9367     Reply to this: 9486
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Xend QPD schematic investigation 

I have looked at the photo of the Xend QPD from elog 9367, as well as the schematic for the board (D990272). 

Things that will need swapping out:

  • Thick film resistors in the signal path need to be changed to thin film.
  • MAX333 needs to be replaced with MAX333A.  The 333 has "ON Resistance" of 140-175 ohms, whereas the 333A has "ON Resistance" of 20-35 ohms.
  • AD797 needs to be replaced by OPA140.  The 797 is a low voltage noise op-amp, but for a diode we want low current noise.  AD797 has 2pA/rtHz at 1kHz, whereas the OP140 has 0.8fA/rtHz at 1kHz (see Zach's elog 8125 re: OPA140).

I have ordered from digikey via techmart 10 each of the MAX333A's and the OPA140's.  (4 per QPD times 2 QPDs plus 2 spares = 10).  Both of these new chips have the same footprint and pinout as the part that they are replacing, so it'll be a fairly easy task.

Next up, I need to make a LISO model for the circuit for one of the quadrants, to see what shape it'll turn out to be.  Part of this will include deciding what resistors and capacitors to put in the OPA140 gain stage. 

Right now, the AD797s say on the schematic that the gain options are different by a factor of 5, but the actual QPD has a different resistor than is on the schematic, and there is also a capacitor in parallel with each resistor, so I need to just pull those out, and pick some values that make sense to me. 

Rana and I have discussed ignoring the 2nd and 3rd gain switching options on each quadrant, as that is getting to be more fine control than we need. 

Other things on the board:

  • The 50 ohm resistors to ground for the "QPD_rtn" have all shorted.  Rana says this is good, so leave it as-is.
  • The positive input to the AD797's all have a 100 ohm resistor to ground, rather than just being connected to ground.  Why is this??

For now, I will probably just work on the QPD head, and not the whitening board.  For now, we can run with 1 stage of whitening, and if we need lower noise, we can revisit the whitening board (including replacing the thick film resistors with thin film).

When thinking about what gains I want on my gain stages, I want to have my full arm power (~700 TRX units) be ~20,000 counts from the ADC.  So, I want my single arm power (1 TRX unit) to be ~30 counts from the ADC.  This is not such a big number, so this may also require more thinking. 

ELOG V3.1.3-